# Boosting Ensembles for Statistics of Tails at Conditionally Optimal Advance Split Times

### v1

@@ 2.2 Estimating conditional and climatological probabilities from boosted ensembles
However, the specific choice of estimator is only auxiliary to our main goal of characterizing the optimal AST. The most important advantage of both estimators over the output from a rare event algorithm, e.g., TEAMS, is an extensible dataset: if the variance is too high, one can always either generate new ancestors by extending the short DNS, or extend the range of ASTs sampled, or enlarge the ensemble at each AST, without discarding the laborious samples already generated. This is unfortunately not the case with an algorithm like TEAMS: because of the random rules by which ancestors are selected and new members generated, a completed run of TEAMS cannot be enlarged while retaining its estimation properties. This results in huge waste during the fine-tuning process of calibrating TEAMS, in contrast to boosted ensembles which can be re-used with different hyperparameter choices.
To emphasize the conditional nature of $A_{j_{n}}$—its possible dependence on the ancestor $n$, due to initial condition-dependent predictability—we refer to $A_{j_{n}}$ as the “conditional advance split time” (CAST), and its optimal value (by $\chi^{2}$ or other criteria) as the “conditional optimal advance split time” (COAST). Our goal is to define the COAST, calculate it given extensive sampling from boosted ensembles, and finally to suggest useful criteria to estimate it when sample size is limited.
@@ 2.3 AST selection criteria
With a data-generating plan and an estimator in place, we return to our central question of interest: how to select the COASTs $\{A_{j_{n}}\}$? There are three natural kinds of criteria.
@@
    Choose a single uniform AST $A_{j_{n}}=A^{\$}$ for all ancestors (\$\\$\$ for “synchronized”). In this case, the CAST is not really “conditional” at all. In Finkel and O’Gorman, (2024), we found the COAST for TEAMS by systematic grid search through candidate ASTs, and found post-hoc an empirical relationship for the COAST: $A^{\$}\approx\overline{t_{3/8}}$, where $\overline{t_{\epsilon}}$ is the average (over the attractor, or equivalently over ancestors) of the time until the ensemble’s root-mean-square distance from the ancestor dispersed to a fraction $\epsilon$ of its saturation value.
@@
    Define an indicator for ensemble dispersion and choose the CAST as the time that the indicator crosses some pre-defined threshold. Specifically, we compute the pattern correlation $\rho$ between spatiotemporal fields $F_{0}$ (from the ancestor) and $F_{m}$ (from the $m$th ensemble member) as
@@
    Unless noted otherwise, $\rho$ will refer to the average of $\rho[F_{0},F_{m}]$ over all members $m=1,\ldots,M$. Pattern correlation is restricted to the range [1,1][-1,1][ - 1 , 1 ] by the Cauchy-Schwarz inequality, and tends to decrease over time from 1 to 0 except for occasional negative values when $F_{0}$ and $F_{1}$ are similar up to translation (but this effect usually disappears when averaging large-enough ensembles). We then choose some threshold $\rho^{\$}\in(0,1)$, and select the corresponding CAST $A_{j_{n}}=A_{n}^{\text{\textcent}}[\rho^{\$}]$—a function of the threshold—as the smallest sampled AST for which the ensemble launched at time $t_{n}^{*}-A_{n}^{\text{\textcent}}$ crosses the threshold by time $t_{n}^{*}$ (¢ for “crossing” the threshold, downward in the case of pattern correlation). Note that the CAST varies with $n$, but the correlation threshold, denoted $\rho^{\$}$, is uniform. Finding the COASTs $A_{n}^{\text{\textcent}}$ then boils down to finding the optimal value of $\rho^{\$}$.
    The $\frac{3}{8}$ rule from Finkel and O’Gorman, (2024), which used euclidean distance $D^{2}[F_{0},F_{m}]=\overline{(F_{0}-F_{m})^{2}}$ as the dispersion indicator, can be approximately restated in terms of pattern correlation:
    $D^{2}$ $=\epsilon^{2}\langle D^{2}\rangle$ $\langle D^{2}\rangle=\text{saturation value of }D^{2}$ (20)
    $\implies\overline{f_{0}^{2}}+\overline{f_{2}^{2}}-2\overline{f_{0}f_{2}}$ $=\epsilon^{2}(\langle{\overline{f_{0}^{2}}}\rangle+\langle{\overline{f_{m}^{2}}}\rangle)$ $\text{Using }\langle{\overline{f_{0}f_{m}}}\rangle=\langle{\overline{f_{0}}}\rangle\langle{\overline{f_{m}}}\rangle=0$ (21)
@@
    $\frac{(1-\epsilon^{2})\langle{\overline{f_{0}^{2}}}\rangle+(1-\epsilon^{2})\langle{\overline{f_{m}^{2}}}\rangle}{\sqrt{\langle{\overline{f_{0}^{2}}\rangle\langle\overline{f_{m}^{2}}}\rangle}}$ $\approx 2\rho(F_{0},F_{m})$ $\text{Approximating }\overline{f_{m}^{2}}\approx\langle{\overline{f_{m}^{2}}}\rangle$ (23)
@@
    In other words, the time until RMSE reaches $\frac{3}{8}$ of its saturation value is roughly equivalent to the time at which pattern correlation drops to $1-(\frac{3}{8})^{2}=0.86$. We do not assume this threshold is optimal, but include it as a reference for comparison. And we stress that the $\frac{3}{8}$ rule implemented in Finkel and O’Gorman, (2024) determines a uniform $A^{\$}$, not a conditional $A^{\text{\textcent}}$, because there averaging was performed over the attractor.
@@
[…]
@@
[…]
      where $(\cdot)_{+}:=\max(\cdot,0)$
@@
      where the levels $r_{k}$ start at $\mu$, and so only the tail part of the conditional CCDF contributes.
@@
    We sometimes write $A^{\pounds}[\text{EI}]$ and $A^{\pounds}[\text{TE}]$ to clarify which functional is being optimized. Where it doesn’t cause confusion, we will also call these COASTs because they are optimizing something, although it is something different than $\chi^{2}$. Our hope is that these two notions of optimality coincide, i.e., by each ancestor separately optimizing EI or TE, the resulting aggregate of distributions (via MoCTail or PoPTail estimators) will minimize $\chi^{2}$-divergence from the true climatological tail.
@@
These criteria are each in turn more complex, but also more theoretically appealing. The correlation-based CASTs $\{A_{n}^{\text{\textcent}}\}_{n=1}^{N_{\text{short}}}$, unlike the synchronized AST $A^{\$}$, can vary with $n$ to respect differences in predictability between different initial conditions, a well-recognized phenomenon in chaotic systems (Maiocchi et al.,, 2024), including the atmosphere (Lucarini and Gritsun,, 2020). Still, both $A^{\$}$ and $A_{n}^{\text{\textcent}}$ require the user to set some arbitrary global threshold , earning them the (pejorative) label “coordinated”, as opposed to the “liberated” $A_{n}^{\pounds}$. The open question is whether optimizing $A_{n}^{\pounds}$ individually for each $n$ will also optimize the accuracy of the unconditional (MoCTail) CCDF against a ground truth.
@@
Main result: Climatological tails are estimated better with perturbed ensembles than with un-perturbed ancestors alone. This holds with few exceptions for all COAST selection rules and across a wide range of target spatial locations. No single selection rule is always superior, nor is either the MoCTail or PoPTail estimators, but a general pattern is that $A^{\$}$ and $A^{\text{\textcent}}$ marginally outperform $A^{\pounds}$[TE], which in turn outperforms $A^{\pounds}$[EI]. The latter two “liberated” criteria, however, have a distinct advantage of needing no arbitrary threshold choices. Furthermore, EI-based estimates, although statistically poor, are useful because they consistently err in a specific direction of over-estimating probabilities (equivalently, severities), giving upper bounds. TE-based estimates strike a reasonable compromise between statistical error and arbitrariness, which is strong enough support that we recommend TE as a generic AST selection rule.
@@
The remainder of the paper demonstrates the theoretical framework above on the QG system. Sect. 3 specifies the dynamical model and its numerical simulation, displays some representative output, defines the target intensity functions of interest, and reports on their basic tail statistics. Sect. 4 specifies the perturbation protocol (i.e., the space $\Omega$ and probability densities $p(\omega)$) and visualizes representative examples of the system’s response, providing motivation for our choices of AST selection criteria. Sect. 6 compares the performances of all proposed AST selection criteria criteria in matching the climatological tail CCDF. Sect. 7 concludes with a summary and outlook on important future lines of work.
@@
Throughout, we present more in-depth results for one select target latitudes just below the domain center, and only summaries for the wider range of target latitudes, which reveals large-scale variations in extreme event predictability and representability across space.
@@ 3 The quasigeostrophic model
The model setup aims to distill some challenges we have encountered with rare event algorithms across the hierarchy. We first recognized the need for advance splitting (or “trying early”) in the context of an aquaplanet GCM (Frierson et al.,, 2007), in which ensembles dispersed too slowly to meaningfully amplify the bursts of rainfall deposited by passing midlatitude cyclones. A minimal surrogate model replicating this challenge was found in Lorenz-96 Lorenz and Emanuel, (1998), which provided a testbed for the first working version of TEAMS and a recognition of an “optimal advance split time” (Finkel and O’Gorman,, 2024). There is a huge gap in model complexity between Lorenz-96 and the GCM (see Table 1), and we wish to test our idea in this middle ground where the target spatial location can have an effect. Lorenz-96, with a one-dimensional domain and homogeneous forcing, is too simple. For this reason, and to make closer contact with physics, we selected the two-layer QG model as a suitable intermediate between Lorenz-96 and the GCM.
Table 1: Three rungs on the model hierarchy
@@ 3.1 Equations of motion and numerical simulation
We implement a version of the QG model combining elements of several classic studies. Our numerical method and friction form follow Haidvogel and Held, (1980), but on a smaller domain as in Panetta, (1993) to contain only 1-2 zonal jets, and with bottom topography in the lower layer as in Thompson, (2010) to fix preferred latitudes for jets while still allowing them to temporarily split, merge, and meander. Thus climate statistics, and hence the COAST itself, can vary with latitude. Further, we augment the system with a passive tracer to represent a key component of precipitation dynamics, following the spirit of Bourlioux and Majda, (2002) and Qi and Majda, (2016, 2018) who used turbulent advection-diffusion as a paradigm for intermittency.
The model equations are as follows, with non-dimensional parameter values listed in Table 2. The horizontal coordinates $(x,y)$ each run from 0 to $L$. The integer-valued vertical coordinate $z$ is an index for the layer (1 for the top and 2 for the bottom). $\psi$ represents the streamfunction minus a background of $-Uy\delta_{z,1}$, where $U$ is an imposed background wind shear. $q$ represents potential vorticity minus a background of $\beta y+h\delta_{z,2}$, due to planetary vorticity gradient and topography. $c$ represents the passive tracer field.
@@
$\Big{[}\partial_{t}+(\partial_{x}\psi)\partial_{y}+(U\delta_{z,1}-\partial_{y}\psi)\partial_{x}\Big{]}c$ =0=0= 0 (28)
@@
For $\psi$, we impose doubly periodic boundary conditions and timestep with a pseudo-spectral method with 64 Fourier modes in each dimension and standard $\frac{2}{3}$-dealiasing (hence, an effective maximum wavenumber of 20). We time-step linear terms with the trapezoid rule (Crank-Nicolson) and nonlinear and topographic terms with a predictor-corrector (Heun’s) method. Meanwhile, boundary conditions on $c$ are periodic in $x$ and Dirichlet in $y$, with values (0,1)(0,1)( 0 , 1 ) at $y=(0,L)$. Together with a first-order upwind monotone finite-volume scheme, this setup guarantees that $0\leq c\leq 1$ everywhere, putting to rest any questions about the boundedness of its probability distribution. Note there is no explicit dissipation for $c$, but the low-order discretization creates some effective diffusivity.
@@
$d=(2\text{ layers})\times(41^{2}\text{ Fourier modes for }\psi+64^{2}\text{ grid cells for $c$})=11554,$ (33)
@@
Table 2: Physical parameters used for the numerical simulation

### v2

@@ 2.3 Evaluating performance
Computational efficiency is another important consideration besides accuracy, as the entire goal of rare event algorithms is to improve efficiency or accuracy (or both) relative to DNS. For a boosting-like rare event algorithm to be useful, its error should decrease faster by perturbing existing ancestors (increasing $M$) than by extending DNS by generating new ancestors (increasing $N$ and not $M$), at least in some range of $N$ that samples the attractor broadly but not exhaustively. However, this paper does not present a complete rare event algorithm per se, in the sense we don’t yet stake our claim on a speedup. Rather, we ask a pre-requisite question: does increasing $M$ decrease the error at all? Clearly boosting can increase the maximum severity, but that could happen in ways that don’t respect the tail CCDF’s shape, e.g., if perturbations tend to maximize the event’s severity while bypassing moderate severities that carry significant statistical weight. We will thus make two comparisons between boosting and DNS: accuracy at fixed $N$, and accuracy at fixed cost (where DNS runs an additional length equal to the cost of simulating descendants, allocating its full budget to “exploration” rather than “exploitation”). Specifically, we approximate the cost of the boosting approach for a given AST $A$ as
$\text{Average boosting cost per ancestor}=M(A+\delta t^{*})+\text{(mean return period)},$ (21)
where $\delta t^{*}$, the “argmax drift” parameter, accounts for the extra time needed to run after the ancestor’s peak to account for changes in peak timing. “Mean return period” is the average time between consecutive independent peaks over the threshold $\mu$, which will be longer than $1/(1-q(\mu))$ because of de-clustering. The dependence on $A$ is a complication, as each AST tried would merit a different-length DNS for cost comparison, and we don’t want to penalize boosting too severely by summing over all ASTs because in practice we would not bother simulating the obviously sub-optimal ASTs. Rather, we optimistically estimate the cost if $A$ is already known. On the other hand, our chosen $M(=21)$ is likely more samples than necessary to fit a satisfactory parametric model, as we have deliberately sampled the perturbation space more generously than we would if chasing a speedup. We simplify the comparison by fixing $A$ to $\frac{1}{2}A_{\text{max}}$ in Eq. (21), which is close to or slightly greater than the optimal values that we found empirically.
We will show that boosting is unambiguously more accurate than DNS when fixing the number of ancestors $N$, and similarly accurate with marginal improvements when fixing cost, though with variation across latitudes and AST criteria. Any fixed-cost performance gains we achieve here (not our main objective) should be viewed as a lower bound for future algorithms, which will benefit from the conceptual insights into AST that we glean presently.
To emphasize the conditional nature of the AST—its possible dependence on the ancestor $n$ due to initial condition-dependent predictability—we refer to $A_{j_{n}}$ as the “conditional advance split time” (CAST), and its optimal value (by $\chi^{2}$ or other criteria) as the “conditionally optimal advance split time” (COAST). Our goal is to define the COAST, calculate it given extensive sampling from boosted ensembles, and finally to suggest useful criteria to estimate it when sample size is limited, as in a real rare event algorithm deployment.
@@ 2.4 AST selection criteria
With a data-generating plan and an estimator in place, we return to our central question of interest: how to select the CASTs $\{A_{j_{n}}\}$? There are three natural kinds of criteria.
@@
    Choose a single uniform AST $A_{j_{n}}=A^{\text{U}}$ for all ancestors (U for “uniform”). In this case, the CAST is not really “conditional” at all. In Finkel and O’Gorman (2024), we found the COAST for TEAMS by systematic grid search through candidate ASTs, and found post-hoc an empirical relationship for the COAST: $A^{\text{U}}\approx\overline{t_{3/8}}$, where $t_{\epsilon}(x_{0})$ is the time until an ensemble dispersing from initial condition $x_{0}$ (each member forced by a different noise realization) reaches a fraction $\epsilon$ of its asymptotic root-mean-squared-error (RMSE), and $\overline{t_{\epsilon}}$ is the average of $t_{\epsilon}(x_{0})$ over different initial conditions $x_{0}$. In Finkel and O’Gorman (2024), we sampled $x_{0}$ from the stationary distribution; here, for computational expediency, we will repurpose the boosting ensembles for estimating $\overline{t_{\epsilon}}$, i.e., sampling $x_{0}$ from pre-peak antecedent conditions.
@@
    Choose the CAST $A_{n}$ separately for each ancestor $n$ such that that an ensemble launched at $t_{n}^{*}-A_{n}$ disperses to a pre-defined threshold at time $t_{n}^{*}$. One could measure dispersal in different ways like RMSE, but here we opt instead for a pattern correlation, defined with respect to spatiotemporal fields $F_{0}$ (from the ancestor) and $F_{m}$ (from the $m$th ensemble member) as
    $\rho[F_{0},F_{m}]$ $:=\frac{\overline{f_{0}f_{m}}}{\sqrt{(\overline{f_{0}^{2}})(\overline{f_{m}^{2}})}}\text{ where }f:=F-\langle F\rangle,\ \ \langle\cdot\rangle=\text{ time-average (climatology), and }\overline{(\cdot)}=\text{ space-average.}$ (22)
    Unless noted otherwise, $\rho$ will refer to the average of $\rho[F_{0},F_{m}]$ over all members $m=1,\ldots,M$. The reason for subtracting time-averages is to fairly weight spatial regions with smaller background $\langle F\rangle$, e.g., poles if $F$ is temperature. Dividing by spatial standard deviations is simply a useful normalization that restricts $\rho$ to the range [1,1][-1,1] by the Cauchy-Schwarz inequality. $\rho$ tends to decrease over time from 1 to 0 except for occasional negative values when $F_{0}$ and $F_{1}$ are similar up to translation (but this effect usually disappears when averaging large-enough ensembles). We then choose some threshold $\rho^{\text{U}}\in(0,1)$, and select the corresponding CAST $A_{j_{n}}=A_{n}^{\text{PC}}[\rho^{\text{U}}]$—a function of the threshold—as the smallest sampled AST $A_{n}$ for which $\rho$ decreases from 1 to $\rho^{\text{PC}}$ between the split time $t_{n}^{*}-A_{n}$ and the peak time $t_{n}^{*}$. (PC stands for for “pattern correlation”). Note that the CAST varies with $n$, but the correlation threshold, denoted $\rho^{\text{U}}$, is uniform. Finding the COASTs $A_{n}^{\text{PC}}$ then boils down to finding the optimal value of $\rho^{\text{U}}$.
    The $\frac{3}{8}$ rule from Finkel and O’Gorman (2024), which used Euclidean distance $D^{2}[F_{0},F_{m}]=\overline{(F_{0}-F_{m})^{2}}=\overline{(f_{0}-f_{m})^{2}}$ as the dispersion indicator, can be approximately restated in terms of pattern correlation:
@@
    $\implies\overline{f_{0}^{2}}+\overline{f_{m}^{2}}-2\overline{f_{0}f_{m}}$ $=\epsilon^{2}(\langle{\overline{f_{0}^{2}}}\rangle+\langle{\overline{f_{m}^{2}}}\rangle)$ $\text{Using }\langle{\overline{f_{0}f_{m}}}\rangle=\langle{\overline{f_{0}}}\rangle\langle{\overline{f_{m}}}\rangle=0$ (24)
@@
    $\frac{(1-\epsilon^{2})\langle{\overline{f_{0}^{2}}}\rangle+(1-\epsilon^{2})\langle{\overline{f_{m}^{2}}}\rangle}{\sqrt{\langle{\overline{f_{0}^{2}}\rangle\langle\overline{f_{m}^{2}}}\rangle}}$ $\approx 2\rho(F_{0},F_{m})$ $\text{Approximating }\overline{f^{2}}\approx\langle{\overline{f^{2}}}\rangle$ (26)
@@
    (The approximation invoked in the second-to-last step, $\overline{f^{2}}\approx\langle\overline{f^{2}}\rangle$, will hold when the spatial region is large enough that global fluctuations in the same direction are unlikely.) This calculation shows that the time until RMSE reaches $\frac{3}{8}$ of its saturation value is roughly equivalent to the time at which pattern correlation drops to $1-(\frac{3}{8})^{2}=0.86$. We do not assume this threshold is optimal, but include it as a reference for comparison. And we stress that the $\frac{3}{8}$ rule implemented in Finkel and O’Gorman (2024) determines a uniform $A^{\text{U}}$, not a conditional $A^{\text{PC}}$, because their averaging was performed over the attractor, whereas here we will use $\rho$ as an initial condition-specific diagnostic.
@@
    Define the CAST as the solution to an optimization problem, where we seek to optimize a functional on the boosted severity distribution that favors both a high mean and high variability of the severity. This would implicitly favor intermediate ASTs, as short-AST ensembles have high mean but low variability while long-AST ensembles will have high variability but low mean (approaching the climatological distribution). We propose and evaluate two such functionals in this paper:
@@
      $\mathbb{E}[(\Delta R^{*})_{+}]=\int_{\Omega}p^{\Omega}(\omega)[R^{*}(\omega)-R^{*}(0)]_{+}\,d\omega,$ (28)
      where $(\cdot)_{+}:=\max(\cdot,0)$ and we recall that $\omega=0$ means no perturbation (i.e., the ancestor)
@@
      where the bin boundaries $r_{k}$ start at $\mu$, and so only the tail part of the conditional CCDF contributes. The thresholded entropy is thus defined based on probability over discrete bins (with the bin boundaries $r_{k}$ set based on quantiles of the ground-truth distribution) and would change if the bins were changed.
@@
    Where it doesn’t cause confusion, we will also call the CASTs $A^{\text{EI}}$ and $A^{\text{TE}}$ themselves COASTs because they optimize something, although it is something different than $\chi^{2}$. We have conjectured that that these two notions of optimality coincide: if each ancestor separately optimizes EI or TE, the resulting aggregate of distributions (via MoCTail or PoPTail estimators) will minimize $\chi^{2}$-divergence from the true climatological tail. Our results will approximately confirm the conjecture in the case of TE.
@@
These criteria are each in turn more complex, but also more theoretically appealing. The correlation-based CASTs $\{A_{n}^{\text{PC}}\}_{n=1}^{N_{\text{short}}}$, unlike the synchronized AST $A^{\text{U}}$, can vary with $n$ to respect differences in predictability between different initial conditions, a well-recognized phenomenon in chaotic systems (Maiocchi et al., 2024), including the atmosphere (Lucarini and Gritsun, 2020). Still, both $A^{\text{U}}$ and $A_{n}^{\text{PC}}$ require the user to set some arbitrary global threshold. The open question is whether optimizing $A_{n}^{\text{EI}}$ or $A_{n}^{\text{TE}}$ individually for each $n$ will also optimize the accuracy of the unconditional (MoCTail) climatological CCDF estimator against the ground truth climatological CCDF from a long DNS.
@@
Main result: Climatological tails are estimated more accurately with perturbed ensembles than with un-perturbed ancestors alone (fixed-$N$ comparison between DNS and boosting). This holds with few exceptions for both MoCTail and PoPTail estimators, for all COAST selection rules, and for all target spatial locations. At fixed cost, boosting and DNS are tied overall, but with some variation across latitudes and the value that cost is fixed to, suggesting that substantial speedups are possible with more highly optimized boosting-like algorithms. No single selection rule is superior across the board. The EI and TE criteria, however, have a distinct advantage of needing no arbitrary threshold choices. TE-based estimates strike a reasonable compromise between statistical error and arbitrariness, which is strong enough support that we recommend TE as a generic AST selection rule.
@@
The remainder of the paper demonstrates the theoretical framework above on the QG system. Sect. 3 specifies the dynamical model and its numerical simulation, displays some representative output, defines the target intensity functions of interest, and reports on their basic tail statistics. Sect. 4 specifies the perturbation protocol (i.e., the space $\Omega$ and probability densities $p^{\Omega}(\omega)$) and visualizes representative examples of the system’s response, providing motivation for our choices of AST selection criteria. Sect. 6 compares the performances of all proposed AST selection criteria criteria in matching the climatological tail CCDF. Sect. 7 concludes with a summary and outlook on important future lines of work.
@@
Throughout, we present more in-depth results for one select target latitudes just south of the domain center, and only summarize for the wider range of target latitudes, which reveals large-scale variations in extreme event predictability and representability across space.
@@ 3 The quasigeostrophic model
The model setup aims to distill some challenges we have encountered with rare event algorithms across the hierarchy. We first recognized the need for advance splitting (or “trying early”) to sample extreme precipitation in an aquaplanet GCM (Finkel and O’Gorman, 2025). A minimal surrogate model replicating this challenge was found in Lorenz-96 Lorenz and Emanuel (1998), which provided a testbed for the first working version of TEAMS and a recognition of an “optimal advance split time” (Finkel and O’Gorman, 2024). There is a huge gap in model complexity between Lorenz-96 and the GCM (see Table 1), and we wish to test our idea in this middle ground where the target spatial location can have an effect. Lorenz-96, with a one-dimensional domain and homogeneous forcing, is too simple. For this reason, and to make closer contact with physics, we selected the two-layer QG model as a suitable intermediate between Lorenz-96 and the GCM.
Table 1: Three rungs on the model hierarchy. Left: the Lorenz-96 system used in Finkel and O’Gorman (2024) has a one-dimensional spatial domain (“longitude”) divided into discrete sites $k=0,\ldots,39$, on which generic meteorological variables $\{x_{k}\}$ evolve in time. Its state space dimension is 40. Right: the aquaplanet model used in Finkel and O’Gorman (2025) has a three-dimensional spatial domain: latitude $\lambda$, longitude $\phi$, and pressure normalized by its surface value, $\sigma=p/p_{s}$. It has six prognostic fields: zonal wind $u$, meridioal wind $v$, temperature $T$, and humidity $q$ vary in all three dimensions, whereas surface pressure $p_{s}$ and precipitation rate $R$ vary only in the horizontal. Center: the 2-layer quasigeostrophic model used in this study has two layers ($z=1,2$) of two dimensions each (longitude $x$, latitude $y$), and two dynamical fields: streamfunction $\psi$ which is discretized spectrally, and tracer concentration $c$ which is discretized on a grid.
@@ 3.1 Equations of motion and numerical simulation
We implement a version of the QG model combining elements of several classic studies. Our numerical method and friction form follow Haidvogel and Held (1980), but on a smaller domain with weaker friction magnitude as in Panetta (1993) to contain only 1-2 more energetic zonal jets. We furthermore add bottom topography in the lower layer as in Thompson (2010) to fix preferred latitudes for jets while still allowing them to temporarily split, merge, and meander. Thus climate statistics, and hence the COAST itself, can vary with latitude. Further, we augment the system with a passive tracer to represent a key component of precipitation dynamics, following the spirit of Bourlioux and Majda (2002) and Qi and Majda (2016, 2018) who used turbulent advection-diffusion as a paradigm for intermittency.
The model equations are as follows, in non-dimensionalized form using the deformation radius $\lambda$ as the length scale and a velocity scale $\mathcal{U}$. To make plain the role of the background shear, we define a non-dimensional wind $U$ as the ratio between the imposed upper-level zonal wind and $\mathcal{U}$. All non-dimensional parameter values are listed in Table 2. The horizontal coordinates $(x,y)$ each run from 0 to $L$. The integer-valued vertical coordinate $z$ is an index for the layer (1 for the top and 2 for the bottom). $\psi$ represents the streamfunction minus a background of $Uy\delta_{z,1}$. $h$ is the bottom topography which is specified to vary sinusoidally with wavenumber 2 in latitude. $q$ represents potential vorticity minus a background of $\beta y+h\delta_{z,2}$, due to planetary vorticity gradient and topography. $c$ represents the passive tracer field.
@@
$\Big[\partial_{t}+(\partial_{x}\psi)\partial_{y}+(U\delta_{z,1}-\partial_{y}\psi)\partial_{x}\Big]c$ $=0$ (31)
@@
For $\psi$, we impose doubly periodic boundary conditions and timestep with a pseudo-spectral method with 64 Fourier modes in each dimension and standard $\frac{2}{3}$-dealiasing (hence, an effective maximum wavenumber of 20). We time-step linear terms with the trapezoid rule (Crank-Nicolson) and nonlinear and topographic terms with a predictor-corrector (Heun’s) method. Meanwhile, boundary conditions on $c$ are periodic in $x$ and Dirichlet in $y$, with values $(0,1)$ at $y=(0,L)$. Together with a first-order upwind monotone finite-volume scheme, this setup guarantees that $0\leq c\leq 1$ everywhere, making clear that its probability distribution has compact support. Note there is no explicit dissipation for $c$, but the low-order discretization creates some effective diffusivity.
@@
[…]
@@
Table 2: Non-dimensional physical parameters used for the numerical simulation, similar to those chosen in Panetta (1993).